OpenStudy (anonymous):

What is the standard form equation of the line passing through (1, –5) with a slope of –1?

OpenStudy (marissalovescats):

y-y1=m(x-x1)

OpenStudy (marissalovescats):

Then put that into y=mx+b

OpenStudy (primeralph):